Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2025 logbook, Contact: F.Wolff-Fabris, A.Galler, H.Sinn  Not logged in ELOG logo
Message ID: 77     Entry time: 31 May 2025, 07:30
Author: Yoonhee Kim 
Group: SPB 
Subject: Shift summary 

X-ray delivery

  • 9.3 keV, 2000 uJ at the beginning, 1600 uJ after 2AM. There was downtime around 15 minutes due to gun issue.
  • FXE used the beam until 23:30.

Optical laser delivery

  • ns laser for particle illumination inside the chamber,
  • another ns laser to visualize the tip of capillary on top of the chamber

Achievements/Observations

  • Measured PSI
  • Hit rate is extremely low. We could see some peaks from AGIPD preview, but it was difficult to see it from the hit finder.

Issues

  • Sample capillary line was clogged twice.
  • From the images of capillary tip and particle scattering (rayleigh / mie scattering), it seems it works fine. But hit rate is too low.

Instrument status

  • AGIPD HV off.
  • AGIPD gate valve closed and pinned.
  • Attenuators set to thickest diamond.
  • Vacuum valves closed.
  • AGIPD breaks enguaged
  • ns Lasers off
  • HIREX is out
ELOG V3.1.4-7c3fd00